# ToolSpectrum : Towards Personalized Tool Utilization for Large Language Models

### v1

@@ Abstract
While integrating external tools into large language models (LLMs) enhances their ability to access real-time information and domain-specific services, existing approaches focus narrowly on functional tool selection following user instructions, overlooking the context-aware personalization in tool selection. This oversight leads to suboptimal user satisfaction and inefficient tool utilization, particularly when overlapping toolsets require nuanced selection based on contextual factors. To bridge this gap, we introduce ToolSpectrum, a benchmark designed to evaluate LLMs’ capabilities in personalized tool utilization. Specifically, we formalize two key dimensions of personalization, user profile and environmental factors, and analyze their individual and synergistic impacts on tool utilization. Through extensive experiments on ToolSpectrum, we demonstrate that personalized tool utilization significantly improves user experience across diverse scenarios. However, even state-of-the-art LLMs exhibit the limited ability to reason jointly about user profiles and environmental factors, often prioritizing one dimension at the expense of the other. Our findings underscore the necessity of context-aware personalization in tool-augmented LLMs and reveal critical limitations for current models. Our data and code are available at https://github.com/Chengziha0/ToolSpectrum.
@@
footnotetext: Equal Contributions. Work done during the first author’s internship at Beihang University.footnotetext: Corresponding Authors
@@ 1 Introduction
  • To best of our knowledge, we are the first to define personalized tool utilization, which proposes a new challenge: current LLMs focus solely on planning tools that fulfill user instructions without considering personalization to select the most suitable tool.
@@ 2.1 Tool Learning Benchmarks
Integrating external tools into LLMs leads to significant advancements in their capabilities, enabling them to perform complex real-world tasks. For instance, tools such as retrievers and calculators enable them to address challenges related to factual accuracy and computation, thereby broadening their potential applications Mialon et al. (2023); Qin et al. (2024). Therefore, evaluating how effectively LLMs utilize these tools becomes a key research focus. Existing benchmarks primarily assess general tool execution performance, including the interaction capabilities of LLMs with tools Qin et al. (2023); Liu et al. (2023b); Huang et al. (2024b); Li et al. (2023), planning and reasoning capabilities with tools Han et al. (2025); Wang et al. (2024b); Huang et al. (2024b); Chen et al. (2024b), as well as the models’ resistance to hallucinations and robustness in tool utilization Zhang et al. (2024a); Huang et al. (2024c); Ning et al. (2024); Zhan et al. (2024); Ye et al. (2024). However, existing benchmarks overlook the presence of the toolset with overlapping functionalities. They typically select the tool complete user’s instruction solely, without considering how user profiles and environmental factors could influence the selection of the most effective tool. To mitigate this gap, we introduce ToolSpectrum, which explicitly considers the impact of user profiles and environmental factors on tool utilization.
@@ 3.2.1 Toolset Collection
In order to cover diverse user instructions and applications, we identify 9 commonly used application domains based on previous studies Guo et al. (2024); Wang et al. (2024b) and app analysis from App Store [^2] https://apps.apple.com/us/charts  [^3] https://play.google.com/store/apps. These domains include shopping, entertainment, travel, delivery, grocery, knowledge, news, health and finance. In each domain, we leverage GPT-4o to generate initial designs for Apps and APIs with similar functions. We then manually curate the outputs, refining them with additional enhancements to ensure high quality. For example, in the shopping domain, users can choose between Amazon or Temu at the APP level to meet the same shopping needs. In the travel domain, users can choose different APIs within the Ctrip [^4] Ctrip is a comprehensive app that offers fast booking for rooms, tickets, and a variety of other travel services. to search for train or flight tickets. We also design corresponding RPs and OPs for each API to meet the reality and diverse user needs following existing works Wang et al. (2024b). The collected toolset is denoted as $\mathcal{T}$, and all Apps and APIs are listed in Table 7.
@@ 3.2.2 Profile and Environment Definition
To model the relationship between user profiles, environments, and personalized tool utilization, we need to create a database that captures both user profiles and environmental factors. In this section, we start with a comprehensive definition of both of these two personalized factors.
@@ 3.2.3 Profile and Environment Collection
For profile, we employ a two-stage quality control process. First, GPT-4o automatically evaluates each profile, assessing Demographic Coherence and Preference Alignment. The scoring range is from 1 to 10. We discard profiles scoring below 8. Second, we manually review the remaining profiles to finalize the profile database $\mathcal{P}$. For environment, since we carefully considered the relationships between various features during the sampling process, no further quality control is needed.
@@ 4.2 Main Results
GPT-4o and DeepSeek-R1 achieve the best overall performance, yet it scores only 0.50 on the $OP$ metric for tasks involving Profile & Environment instructions, highlighting the difficulties posed by such tasks. Models like DeepSeek-V3 and QwQ-32B perform similarly to GPT-4o on certain metrics, reducing the gap between open and closed-source models. Importantly, scaling up model size does not significantly improve personalized tool utilization. For instance, Qwen2.5-72B-Instruct showes little improvement over Qwen2.5-32B-Instruct, suggesting that simply increasing model size may not be effective for complex tasks.
@@ 5 Analysis
In this section, we conduct a comprehensive analysis to answer three research questions RQ1: Is personalization truly necessary when LLMs utilize tools? (Sec 5.1) RQ2: What are the differences in personalized capabilities of LLMs in different domains? (Sec 5.2) RQ3: What are the bottlenecks of LLMs in personalized tool utilization, (Sec 5.3) and can prompt strategy to alleviate it? (Sec B.2 and Sec B.3)
@@ 5.1 The Importance of Personalization in Tool Utilization
This section investigates how integrating personalized factors enhances the effectiveness of tool utilization. Specifically, we compare the results of tool utilization with and without personalized information. We then use GPT-4o to evaluate which output from the two approaches more effectively aligns with user’s contexts, with the prompt template shown in Figure 15. To ensure robustness, we supplement the automated evaluation with human evaluation. We randomly sampled 100 outputs for manual evaluation and calculate the Kappa coefficient between GPT-4o’s and human scores. The Kappa coefficient is 0.85, indicating a very high level of agreement between the GPT-4o’s assessments and human evaluations. As illustrated in Figure 3, two key conclusions can be drawn: (1) The incorporation of personalized factors improves the effectiveness of tool utilization. In general, both GPT-3.5 and GPT-4o show higher effectiveness when personalization is applied. (2) The effectiveness of personalization increases with model strength. While both models benefit from personalization, GPT-4o shows a more significant improvement, as more advanced models better understand the relationship between personalized features and their corresponding parameters.
@@ 5.2 Analysis of Model Performance Across Various Data Types in ToolSpectrum
To better understand the effects of personalization across different domains, we analyze the average performance of four metrics (i.e., APP, API, RP, and OP). Figure 4 shows the final results. On the one hand, when considering a single personalized factor, the model’s performance varies significantly across domains. The worst performance occurs in the News domain, whether the model considers only user profile or environment. We attribute this to the fact that user preferences for News are primarily derived from natural language descriptions (personality) in ToolSpectrum. Compared to features represented in key-value pairs, such descriptions are generally more challenging to interpret.
@@ 5.3 Error Analysis
In the Profile & Environment data type, the model fails to interpret the domain policy correctly. Despite some instructions violating the policy, the model still generates and returns calling results that do not meet expectations, suggesting a lack of understanding or failure to comply with given constraints.
@@ 6 Conclusion
This paper introduces personalized tool utilization, which considers functionally similar tools and user-specific factors to optimize tool utilization in real-world scenarios. Specifically, we define two critical factors: user profile and environment and present ToolSpectrum, a benchmark designed to evaluate personalized tool utilization. Through extensive experiments conducted on ToolSpectrum, we demonstrate that incorporating personalized tool utilization significantly enhances its effectiveness. However, current LLMs face challenges in performing effectively on this new task.
@@ Appendix C Prompt Templates
Prompt Template for Generating Instruction I am creating a dataset for tool calls, and I will input some keywords for you. Please help me output corresponding user commands based on the keywords. requirement: 1. The input includes an API description and keywords. The API description is an introduction to the current application scenario; Keywords are a dictionary containing some keywords, and you need to concatenate them into instructions that fit the user’s usage habits. 2. The generated instruction is in the first person. 3. The generated instructions must include these keywords and cannot be modified in any way. 4. Do not output any irrelevant content, instructions should not contain any irrelevant symbols, and only natural language is allowed. 5. Do not include keys in the keyword dictionary, as the generated instructions can only contain values. 6. Do not explicitly express the content of the API Description. ### Input: ##API Description {api description} ## Keywords: {keywords} ### Output:
@@
Prompt Template for Scoring Tool-call Result I am currently evaluating the quality of tool-calling datasets. I will provide you with user instructions, user profile, external environment, and tool calling results. Please comprehensively consider these factors and tell me whether the tool calling results meet expectations. Requirements: 1. Please judge whether the tool calling results meet expectations based on user instructions, user profile, external environment, and tool calling results. 2. Please rate the following three aspects separately, ranging from 1 to 10 points. The output format is separated by commas. 3. You should evaluate whether the tool calling results meet expectations from three aspects: - Whether the tool calling results can solve the user’s needs; - Whether the tool calling results match the user’s profile; - Whether the tool calling results align with the external environment. 4. Do not output any irrelevant content, only output the answer. Below are the user instructions, user profile, external environment, and tool calling results I’m inputting: ### Input ## User Instruction {instruction} ## User profile {profile} ## Environment {environment} ## Tool Calling Result {tool call result} ### Output

### v2

@@ Abstract
While integrating external tools into large language models (LLMs) enhances their ability to access real-time information and domain-specific services, existing approaches focus narrowly on functional tool selection following user instructions, overlooking the context-aware personalization in tool selection. This oversight leads to suboptimal user satisfaction and inefficient tool utilization, particularly when overlapping toolsets require nuanced selection based on contextual factors. To bridge this gap, we introduce ToolSpectrum, a benchmark designed to evaluate LLMs’ capabilities in personalized tool utilization. Specifically, we formalize two key dimensions of personalization, user profile and environmental factors, and analyze their individual and synergistic impacts on tool utilization. Through extensive experiments on ToolSpectrum, we demonstrate that personalized tool utilization significantly improves user experience across diverse scenarios. However, even state-of-the-art LLMs exhibit the limited ability to reason jointly about user profiles and environmental factors, often prioritizing one dimension at the expense of the other. Our findings underscore the necessity of context-aware personalization in tool-augmented LLMs and reveal critical limitations in current models. Our data and code are available at https://github.com/BUAA-IRIP-LLM/ToolSpectrum.
@@
footnotetext: Equal Contributions. Work done during the first author’s internship at Beihang University.footnotetext: Corresponding Author
@@ 1 Introduction
  • To the best of our knowledge, we are the first to define personalized tool utilization, which proposes a new challenge: current LLMs focus solely on planning tools that fulfill user instructions without considering personalization to select the most suitable tool.
@@ 2.1 Tool Learning Benchmarks
Integrating external tools into LLMs leads to significant advancements in their capabilities, enabling them to perform complex real-world tasks. For instance, tools such as retrievers and calculators enable them to address challenges related to factual accuracy and computation, thereby broadening their potential applications Mialon et al. (2023); Qin et al. (2024). Therefore, evaluating how effectively LLMs utilize these tools becomes a key research focus. Existing benchmarks primarily assess general tool execution performance, including the interaction capabilities of LLMs with tools Qin et al. (2023); Liu et al. (2023b); Huang et al. (2024b); Li et al. (2023), planning and reasoning capabilities with tools Han et al. (2025); Wang et al. (2024b); Huang et al. (2024b); Chen et al. (2024b), as well as the models’ resistance to hallucinations and robustness in tool utilization Zhang et al. (2024a); Huang et al. (2024c); Ning et al. (2024); Zhan et al. (2024); Ye et al. (2024). However, existing benchmarks overlook the presence of the toolset with overlapping functionalities. They typically select the tool to complete the user’s instruction solely, without considering how user profiles and environmental factors could influence the selection of the most effective tool. To mitigate this gap, we introduce ToolSpectrum, which explicitly considers the impact of user profiles and environmental factors on tool utilization.
@@ 3.2.1 Toolset Collection
To cover diverse user instructions and applications, we identify 9 commonly used application domains based on previous studies Guo et al. (2024); Wang et al. (2024b) and app analysis from App Store [^2] https://apps.apple.com/us/charts  [^3] https://play.google.com/store/apps. These domains include shopping, entertainment, travel, delivery, grocery, knowledge, news, health and finance. In each domain, we leverage GPT-4o to generate initial designs for Apps and APIs with similar functions. We then manually curate the outputs, refining them with additional enhancements to ensure high quality. For example, in the shopping domain, users can choose between Amazon or Temu at the APP level to meet the same shopping needs. In the travel domain, users can choose different APIs within the Ctrip [^4] Ctrip is a comprehensive app that offers fast booking for rooms, tickets, and a variety of other travel services. to search for train or flight tickets. We also design corresponding RPs and OPs for each API to meet the reality and diverse user needs, following existing works Wang et al. (2024b). The collected toolset is denoted as $\mathcal{T}$, and all Apps and APIs are listed in Table 7.
@@ 3.2.2 Profile and Environment Definition
To model the relationship between user profiles, environments, and personalized tool utilization, we need to create a database that captures both user profiles and environmental factors. In this section, we start with a comprehensive definition of both of these personalized factors.
@@ 3.2.3 Profile and Environment Collection
For the profile, we employ a two-stage quality control process. First, GPT-4o automatically evaluates each profile, assessing Demographic Coherence and Preference Alignment. The scoring range is from 1 to 10. We discard profiles scoring below 8. Second, we manually review the remaining profiles to finalize the profile database $\mathcal{P}$. For the environment, since we carefully considered the relationships between various features during the sampling process, no further quality control is needed.
@@ 4.2 Main Results
GPT-4o and DeepSeek-R1 achieve the best overall performance, yet it scores only 0.50 on the $OP$ metric for tasks involving Profile & Environment instructions, highlighting the difficulties posed by such tasks. Models like DeepSeek-V3 and QwQ-32B perform similarly to GPT-4o on certain metrics, reducing the gap between open and closed-source models. Importantly, scaling up model size does not significantly improve personalized tool utilization. For instance, Qwen2.5-72B-Instruct shows little improvement over Qwen2.5-32B-Instruct, suggesting that simply increasing model size may not be effective for complex tasks.
@@ 5 Analysis
In this section, we conduct a comprehensive analysis to answer three research questions RQ1: Is personalization truly necessary when LLMs utilize tools? (Sec 5.1) RQ2: What are the differences in personalized capabilities of LLMs in different domains? (Sec 5.2) RQ3: What are the bottlenecks of LLMs in personalized tool utilization, (Sec 5.3) and can a prompt strategy alleviate it? (Sec B.2 and Sec B.3)
@@ 5.1 The Importance of Personalization in Tool Utilization
This section investigates how integrating personalized factors enhances the effectiveness of tool utilization. Specifically, we compare the results of tool utilization with and without personalized information. We then use GPT-4o to evaluate which output from the two approaches more effectively aligns with the user’s contexts, with the prompt template shown in Figure 15. To ensure robustness, we supplement the automated evaluation with human evaluation. We randomly sample 100 outputs for manual evaluation and calculate the Kappa coefficient between GPT-4o’s and human scores. The Kappa coefficient is 0.85, indicating a very high level of agreement between the GPT-4o’s assessments and human evaluations. As illustrated in Figure 3, two key conclusions can be drawn: (1) The incorporation of personalized factors improves the effectiveness of tool utilization. In general, both GPT-3.5 and GPT-4o show higher effectiveness when personalization is applied. (2) The effectiveness of personalization increases with model strength. While both models benefit from personalization, GPT-4o shows a more significant improvement, as more advanced models better understand the relationship between personalized features and their corresponding parameters.
@@ 5.2 Analysis of Model Performance Across Various Data Types in ToolSpectrum
To better understand the effects of personalization across different domains, we analyze the average performance of four metrics (i.e., APP, API, RP, and OP). Figure 4 shows the final results. On the one hand, when considering a single personalized factor, the model’s performance varies significantly across domains. The worst performance occurs in the News domain, whether the model considers only the user profile or the environment. We attribute this to the fact that user preferences for News are primarily derived from natural language descriptions (personality) in ToolSpectrum. Compared to features represented in key-value pairs, such descriptions are generally more challenging to interpret.
@@ 5.3 Error Analysis
In the Profile & Environment data type, the model fails to interpret the domain policy correctly. Despite some instructions violating the policy, the model still generates and returns calling results that do not meet expectations, suggesting a lack of understanding or failure to comply with the given constraints.
@@ 6 Conclusion
This paper introduces personalized tool utilization, which considers functionally similar tools and user-specific factors to optimize tool utilization in real-world scenarios. Specifically, we define two critical factors: user profile and environment, and present ToolSpectrum, a benchmark designed to evaluate personalized tool utilization. Through extensive experiments conducted on ToolSpectrum, we demonstrate that incorporating personalized tool utilization significantly enhances its effectiveness. However, current LLMs face challenges in performing effectively on this new task.
@@ Appendix C Prompt Templates
Prompt Template for Generating Instruction I am creating a dataset for tool calls, and I will input some keywords for you. Please help me output the corresponding user commands based on the keywords. requirement: 1. The input includes an API description and keywords. The API description is an introduction to the current application scenario; Keywords are a dictionary containing some keywords, and you need to concatenate them into instructions that fit the user’s usage habits. 2. The generated instruction is in the first person. 3. The generated instructions must include these keywords and cannot be modified in any way. 4. Do not output any irrelevant content, instructions should not contain any irrelevant symbols, and only natural language is allowed. 5. Do not include keys in the keyword dictionary, as the generated instructions can only contain values. 6. Do not explicitly express the content of the API Description. ### Input: ##API Description {api description} ## Keywords: {keywords} ### Output:
@@
Prompt Template for Scoring Tool-call Result I am currently evaluating the quality of tool-calling datasets. I will provide you with user instructions, user profile, external environment, and tool calling results. Please comprehensively consider these factors and tell me whether the tool’s calling results meet expectations. Requirements: 1. Please judge whether the tool calling results meet expectations based on user instructions, user profile, external environment, and tool calling results. 2. Please rate the following three aspects separately, ranging from 1 to 10 points. The output format is separated by commas. 3. You should evaluate whether the tool calling results meet expectations from three aspects: - Whether the tool calling results can solve the user’s needs; - Whether the tool calling results match the user’s profile; - Whether the tool calling results align with the external environment. 4. Do not output any irrelevant content, only output the answer. Below are the user instructions, user profile, external environment, and tool calling results I’m inputting: ### Input ## User Instruction {instruction} ## User profile {profile} ## Environment {environment} ## Tool Calling Result {tool call result} ### Output